# A predictive model for bubble–particle collisions in turbulence

### v1

@@ 4.4 Collision kernel
We now insert the model $f(w_{b})$ based on (16) and $\sigma^{(BC)}$ into (11) and compare the collision kernel predictions with the values from simulations. Figures 5(ab) show that the predicted and measured values of the nondimensionalised collision kernels agree reasonably well with each other for $St_{b}=2.8$ and $6.3$ over the entire tested range of $St_{p}$ when $1/Fr\geq 2$. Consistent with the simulations, our model captures the general trend that $\Gamma\tau_{\eta}/r_{c}^{3}$ increases with both $St_{p}$ and $1/Fr$. We scrutinise this agreement by plotting the ratio of the collision kernel from simulation $\Gamma^{sim}$ to that obtained from the full model $\Gamma^{model}$ in figures 5(cd). The difference between $\Gamma^{sim}$ and $\Gamma^{model}$ is generally less than 30% when $1/Fr\geq 2$ , except for $St_{p}\sim 0.1$ where the discrepancy becomes larger. $St_{p}\sim 0.1$ corresponds to the transition regime where the interception and inertial mechanisms both play a comparable role, such that $E_{c}$ increases rapidly with $St_{p}^{\prime}$ (Jiang and Krug, 2025). Since $St_{p}^{\prime}$ is based on $w_{b}$, the shape of the entire distribution $f(w_{b})$, not just its mean value, can influence the predicted collision kernel for intermediate values of $St_{p}$, unlike at extreme values of $St_{p}$ where the effect is less pronounced. The discrepancy between $\Gamma^{model}$ and $\Gamma^{sim}$ around $St_{p}=0.1$ is hence due to the fact that our expression of $E_{c}$ (7) with the fitted parameters does not completely capture the interception-to-inertial transition of the collision efficiency even in a uniform background flow, as well as due to the difference between $f(w_{b})$ obtained from the model and the simulations.
@@
To put the accuracy of our model into perspective, we include the prediction by Kostoglou et al. (2020a), $\Gamma^{(Kos)}$, in figures 5(ab). Although their model was developed for particles with small $St_{p}$, it still overestimates the collision kernel even for tracer particles, with $\Gamma^{(Kos)}$ showing a discrepancy at least ten times larger than that of our model. In view of the excellent agreement of our model prediction with simulation data, we use our model prediction as a benchmark and plot $\Gamma^{(Kos)}/\Gamma^{model}$ for $St_{p}=0$ in figure 5(e). The figure shows that the overprediction persists for all $1/Fr\in[1,10]$. We note that in contrast with our model, Kostoglou et al. (2020a) considers the relative motion driven by small-scale shear. This is done by adding to the (radial) collision velocity a contribution $u_{pt}$ that is uniform over all positions on the sphere with radius $r_{c}$ ($u_{pt}=r_{p}/2\cdot\langle|\partial u_{i}\partial x_{i}|\rangle=r_{p}/2\cdot\sqrt{2/(15\pi)}\sqrt{\varepsilon/\nu}$ in (1.2), where the factor 1/2 is to account only for the influx). Remarkably, despite not including this effect in our model, our predictions agree well with the data for tracer particles, suggesting that this shear mechanism likely plays a relative minor role in determining the collision kernel at the present parameters. We remove this mechanism from $\Gamma^{(Kos)}$ to obtain $\Gamma^{(Kos)}_{nS}$. As figure 5(e) shows, $\Gamma^{(Kos)}_{nS}<\Gamma^{(Kos)}$ as expected, yet $\Gamma^{(Kos)}_{nS}$ still significantly and consistently overpredicts the collision rate. We therefore consider another key difference with Kostoglou et al. (2020a): the position of collisions. $\Gamma^{(Kos)}_{nS}$ includes a significant number of collisions in the bubble wake, which is attributed to the inward branches of the bubble wake vortex pair (and is enhanced by the now-neglected small-scale fluid shear effect). Such collisions are very rarely observed in the simulations in this study as well as those in Jiang and Krug (2025) and Tiedemann and Fröhlich (2025) because most of the incident particles collide with the upstream hemisphere of the bubble and are subsequently reseeded, such that the particle concentration is especially low in the bubble wake. This collision scheme corresponds to the scenario where particles attach to bubbles after collision, meaning similar behaviour might be expected in reality. Removing these collisions in the wake (i.e. setting $\theta_{d}=\pi$ in (39)) gives $\Gamma^{(Kos)}_{nSnW}$ in figure 5(e). Since the values of $\Gamma^{(Kos)}_{nSnW}$ are still larger than those of $\Gamma^{model}$, collisions in the bubble wake alone cannot fully account for the overprediction of the model by Kostoglou et al. (2020a).
@@
where $X$, $Y$, and $F$ are defined in (44), $\theta$ is the polar angle with $\theta=0$ pointing upstream, $u_{\infty}$ is the freestream velocity, and positive values of $u_{r}^{(Kos)}$ correspond to flow away from the bubble. This expression is directly tested in figure 5(f) for the case of a bubble in a uniform incident flow with $Re_{b}^{\infty}=2r_{b}u_{\infty}/\nu=120$ and $r_{p}/r_{b}=1/30$ (see A for simulation details). $Re_{b}^{\infty}=120$ is chosen as it is the approximate value of $\langle Re_{b}\rangle$ for the turbulent cases with $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,4)$. Figure 5(f) shows that $-u_{r}^{(Kos)}(r_{c})$ slightly overpredicts the radial fluid velocity in the bubble wake, where the actual number of collisions is low due to the low particle concentration as discussed above. The radial fluid velocity on the upstream ($\theta<$$$ 60  °) side of the bubble, where most of the collisions occur in the simulations, shows significant deviation between the fit and the data. Surprisingly, the values of $-u_{r}^{(Kos)}/u_{\infty}$ are much lower in this region compared to the simulation results. This implies that the overprediction of $\Gamma^{(Kos)}_{nSnW}$, at least for $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,4)$ at the present value of $Re_{\lambda}$, is driven by a significant overprediction of the freestream bubble slip velocity in turbulence, which we partly attributed further to the modelling of the vertical component in their case. There, the mean vertical slip velocity results from a simple force balance in still fluid without considering the reduction in turbulence as discussed in §2.1.
@@
Two main conclusions can be drawn from the above analysis. First, $\Gamma^{(Kos)}$ consistently overestimates the collision kernel in the present parameters, partly because of collisions in the wake which become more frequent in the presence of small-scale shear. However, even when these two effects are excluded, $\Gamma^{(Kos)}_{nSnW}$ still overpredicts the collision kernel since the freestream slip velocity and the distorted flow field are not accurately modelled. For $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,4)$ at the present value of $Re_{\lambda}$, the freestream bubble slip velocity is overpredicted, though this is partially compensated by the poor agreement between the fit of the distorted flow (35) and the data. Therefore, caution is advised when using (35), as its accuracy may be limited under certain flow conditions. Second, our model, in spite of its simplicity, captures the essential phenomena and produces predictions that are much closer to the simulated values compared to the those by Kostoglou et al. (2020a), even in the tracer particle limit.
@@ 5.1 Model validity for the cases with lower bubble Stokes number
where $\rho_{w}(\tau)=\langle w_{b}(t)w_{b}(t+\tau)\rangle_{t}/\langle w_{b}^{2}(t)\rangle_{t}$ is the autocorrelation function of the bubble slip velocity, $\langle\cdots\rangle_{t}$ denotes averaging over time, and $\tau_{\infty}$ is defined as the value of $\tau$ where $\rho_{w}(\tau)$ first decays to 0.01. $T_{L}^{w}/\tau_{f}\gtrsim 1$ would suggest that the frozen turbulence assumption is valid and the model can be applied. The measured values of $T_{L}^{w}/\tau_{f}$ is shown in figure 6. $T_{L}/\tau_{f}$ generally increases with $1/Fr$ and is far less sensitive to $St_{b}$, such that for all $St_{b}$ considered here the frozen turbulence assumption can be applied when $1/Fr\geq 4$. For $St_{b}=0.5$ and 1, we expect that modelling the slip velocity p.d.f. $f(w_{b})$ by (16) and (28) might slightly underpredict the collision kernel, since the mean value of the slip velocity magnitude is underpredicted as shown in figures 4(ab). This discrepancy in $f(w_{b})$ for $1/Fr\geq 4$ is not due to using the mean vertical rise velocity instead of the magnitude of the bubble rise velocity in the bubble drag correction factor $f_{b}$ in (28), because the bubble velocity at high $1/Fr$ is dominated by its vertical component. As evidence, we plot $f(w_{b})$ with $\sigma_{i}^{(BC)}$ determined using the average $f_{b}$ obtained from the simulations as dashed green lines in figure 4. Only at small $1/Fr$ does using the actual value of $f_{b}$ noticeably improve the quality of the predicted slip velocity p.d.f.. If desired for these cases, a more accurate estimate of $\sigma_{i}^{(BC)}$ can be obtained a priori by solving (28) iteratively using $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ as an initial guess. We additionally note that particles with $St_{p}\sim 1$ preferentially sample flow regions in turbulence which can reduce $\Gamma$ by up to 20% (Jiang and Krug, 2025). Similarly preferential sampling occurs for bubbles with $St_{b}\sim 1$ but they cluster at different regions, which means that bubbles and particles segregate spatially when $(St_{b},St_{p})\sim(1,1)$ unless $1/Fr\gtrsim 10$ where segregation becomes negligible under strong gravity (Chan et al., 2025). This segregation effect is not included in our model and reduces the collision kernel, hence for $(St_{b},St_{p})\sim(1,1)$ at intermediate values of $1/Fr$ the predicted value of the collision kernel may be closer to the actual value than anticipated.
@@ 5.2.1 Relevant parameter space
Depending on the specific value within the broad range of possible dissipation rates $\varepsilon$, the maximum stable bubble size varies because bubbles tend to break up when the local fluid shear becomes sufficiently large (Kolmogorov, 1949; Hinze, 1955). Quantitatively, this behaviour is captured by a critical bubble Weber number $We_{ISR}=2.13\rho_{f}(2r_{b}\varepsilon)^{2/3}(2r_{b})/\gamma$ beyond which a bubble cannot remain stable (Masuk et al., 2021). Here, the velocity scaling in the inertial subrange is used and $\gamma=73\,$\mathrm{m}\mathrm{N}\mathrm{/}\mathrm{m}$$ is the surface tension of water. The nominal value for $We_{crit}$ reported in the literature vary (Hinze, 1955; Risso and Fabre, 1998), which is indicated in figure 7(a) along with the values of $We_{ISR}$ for the combinations of $r_{b}$ and $\varepsilon$ investigated here. Figure 7(a) shows that while most of the combinations can stably exist when accounting for the spread in $We_{crit}$, bubbles with $r_{b}=2\,$\mathrm{m}\mathrm{m}$$ would break up when $\varepsilon=100\,$\mathrm{W}\mathrm{/}\mathrm{k}\mathrm{g}$$. This is further consistent with the very high average bubble Reynolds number, $\langle Re_{b}\rangle=2r_{b}\langle|\mathbf{w_{b}}|\rangle/\nu$, estimated by our model for this case. As shown in figure 7(b), $\langle Re_{b}\rangle$ exceeds 6000. Based on these results, we will not evaluate our model predictions for the $(r_{b},\varepsilon)=(2\,$\mathrm{m}\mathrm{m}$,100\,$\mathrm{W}\mathrm{/}\mathrm{k}\mathrm{g}$)$ case.

### v2

@@ 4.4 Collision kernel with non-settling particles
To test the effects of particle inertia and gravity independently, we initially consider the collision kernel for the cases where the particle settling velocity is set to zero. We insert the model $f(w_{b})$ based on (14) and $\sigma^{(BC)}$ into (3) and compare the collision kernel predictions with the values from simulations. Figures 5(ab) show that the predicted and measured values of the nondimensionalised collision kernels agree reasonably well with each other for $St_{b}=2.8$ and $6.3$ over the entire tested range of $St_{p}$ when $1/Fr\geq 2$. Consistent with the simulations, our model captures the general trend that $\Gamma\tau_{\eta}/r_{c}^{3}$ increases with both $St_{p}$ and $1/Fr$. We scrutinise this agreement by plotting the ratio of the collision kernel from simulation $\Gamma^{sim}$ to that obtained from the full model $\Gamma^{model}$ in figures 5(cd). The difference between $\Gamma^{sim}$ and $\Gamma^{model}$ is generally less than 30% when $1/Fr\geq 2$ , except for $St_{p}\sim 0.1$ where the discrepancy becomes larger. $St_{p}\sim 0.1$ corresponds to the transition regime where the interception and inertial mechanisms both play a comparable role, such that $E_{c}$ increases rapidly with $St_{p}^{\prime}$ (Jiang and Krug, 2025). Since $St_{p}^{\prime}$ is based on $w_{b}$, the shape of the entire distribution $f(w_{b})$, not just its mean value, can influence the predicted collision kernel for intermediate values of $St_{p}$, unlike at extreme values of $St_{p}$ where the effect is less pronounced. The discrepancy between $\Gamma^{model}$ and $\Gamma^{sim}$ around $St_{p}=0.1$ is hence due to the fact that our expression of $E_{c}$, i.e. (9), with the fitted parameters does not completely capture the interception-to-inertial transition of the collision efficiency even in a uniform background flow, as well as due to the difference between $f(w_{b})$ obtained from the model and the simulations.
@@
To put the accuracy of our model into perspective, we consider the model of Kostoglou et al. (2020a), which requires the same inputs as ours, and include their prediction, $\Gamma^{(Kos)}$, in figures 5(ab). Although their model was developed for particles with small $St_{p}$, it still overestimates the collision kernel even for tracer particles, with $\Gamma^{(Kos)}$ showing a discrepancy at least ten times larger than that of our model. In view of the excellent agreement of our model prediction with simulation data, we use our model prediction as a benchmark and plot $\Gamma^{(Kos)}/\Gamma^{model}$ for $St_{p}=0$ in figure 5(e). The figure shows that the overprediction persists for all $1/Fr\in[1,10]$. We note that in contrast with our model, Kostoglou et al. (2020a) considers the relative motion driven by small-scale shear. This is done by adding to the (radial) collision velocity a contribution $v_{pt}$ that is uniform over all positions on the sphere with radius $r_{c}$ ($v_{pt}=r_{p}/2\cdot\langle|\partial u_{i}\partial x_{i}|\rangle=r_{p}/2\cdot\sqrt{2/(15\pi)}\sqrt{\varepsilon/\nu}$ in (B), where the factor 1/2 is to account only for the influx). Remarkably, despite not including this effect in our model, our predictions agree well with the data for tracer particles, suggesting that this shear mechanism likely plays a relative minor role in determining the collision kernel at the present parameters. We remove this mechanism from $\Gamma^{(Kos)}$ to obtain $\Gamma^{(Kos)}_{nS}$. As figure 5(e) shows, $\Gamma^{(Kos)}_{nS}<\Gamma^{(Kos)}$ as expected, yet $\Gamma^{(Kos)}_{nS}$ still significantly and consistently overpredicts the collision rate. We therefore consider another key difference with Kostoglou et al. (2020a): the position of collisions. $\Gamma^{(Kos)}_{nS}$ includes a significant number of collisions in the bubble wake, which is attributed to the inward branches of the bubble wake vortex pair (and is enhanced by the now-neglected small-scale fluid shear effect). Such collisions are very rarely observed in the simulations in this study as well as those in Jiang and Krug (2025) and Tiedemann and Fröhlich (2025) because most of the incident particles collide with the upstream hemisphere of the bubble and are subsequently reseeded, such that the particle concentration is especially low in the bubble wake. This collision scheme corresponds to the scenario where particles attach to bubbles after collision, meaning similar behaviour might be expected in reality. Removing these collisions in the wake (i.e. setting $\theta_{d}=\pi$ in (39)) gives $\Gamma^{(Kos)}_{nSnW}$ in figure 5(e). Since the values of $\Gamma^{(Kos)}_{nSnW}$ are still larger than those of $\Gamma^{model}$, collisions in the bubble wake alone cannot fully account for the overprediction of the model by Kostoglou et al. (2020a).
@@
[…]
@@
Two main conclusions can be drawn from the above analysis. First, $\Gamma^{(Kos)}$ consistently overestimates the collision kernel in the present parameters, partly because of collisions in the wake which become more frequent in the presence of small-scale shear. However, even when these two effects are excluded, $\Gamma^{(Kos)}_{nSnW}$ still overpredicts the collision kernel since the freestream slip velocity and the distorted flow field are not accurately modelled. For $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,4)$ at the present value of $Re_{\lambda}$, the freestream bubble slip velocity is overpredicted, though this is partially compensated by the poor agreement between the fit of the distorted flow (33) and the data. Therefore, caution is advised when using (33), as its accuracy may be limited under certain flow conditions. Second, our model, in spite of its simplicity, captures the essential phenomena and produces predictions that are much closer to the simulated values compared to those by Kostoglou et al. (2020a), even in the tracer particle limit.
4.5 Collision kernel with settling particles
Now, we additionally consider the effect of particle gravity on the bubble–particle collision rate. As expected, the simulation results for the collision kernel in figure 6 (a–b) generally increase when particle settling is included, with the effect being most pronounced at large $St_{p}$. Generally, the simulation data is well captured by the model prediction taking gravity into account (plotted as solid lines) for both $St_{b}=2.8$ (in figure 6 (a)) and $St_{b}=6.3$ (figure 6 (b)). At small particle Stokes numbers, the settling is relatively weak, hence the impact of gravity-induced settling on collisions is limited in this regime. The settling velocity increases with increasing $St_{p}$. As a result, the collision kernel grows even more rapidly with increasing $St_{p}$ compared to the case without settling, implying that particle settling becomes the dominant mechanism at large $St_{p}$. This is particularly evident from the modelled collision kernel when nondimensionalised with $r_{c}^{3}/\tau_{\eta}$, which continues to increase beyond $St_{p}\approx 1$ when particle settling is included. We also point out that the difference between settling and non-settling particles is more pronounced at the lower value of $St_{b}$ considered here. This is because the ratio between the particle settling velocity and the bubble rising velocity becomes larger at lower $St_{b}$ (the settling velocity is solely determined by $St_{p}$). Furthermore, we note that the model predictions show better agreement with the simulation data at lower $St_{p}$ when $1/Fr>2$, which is consistent with behaviour observed in the non-settling case.
To further assess the level of agreement, we plot in figure 6 (c-d) the ratio of the collision kernel obtained from simulations to that from the model. We note that Kostoglou et al. (2020a) additionally tried to incorporate the transient misalignment between the bubble slip velocity vector and gravity through a correction coefficient $\mathfrak{f}$ (see details in B). However, adopting this coefficient did not improve the prediction and therefore this secondary effect is neglected in the present model for simplicity. Despite this simplification, the model prediction shows good agreement for cases of $1/Fr<2$, with deviations generally within 20%, which verifies the correction coefficient has only a minor influence on the overall prediction. Furthermore, the results indicate that the discrepancy between $\Gamma^{sim}$ and $\Gamma^{model}$ is reduced in the intermediate $St_{p}$ regime compared with the non-settling case (cf. figure 6(c,d)). This is because particle settling exerts a significant influence in the intermediate $St_{p}$ range. Consequently, the deficiency associated with the modelling of $E_{c}^{(in)}$ in the intermediate $St_{p}^{\prime}$ regime, as discussed in §4.4, becomes less significant, yielding improved agreement between the simulations and model predictions. This finding further underscores the important role of particle settling in regulating the collision process.
Figure 6: (a) The actual and predicted collision kernels at $St_{b}=2.8$ and (b) $St_{b}=6.3$ with settling particles. The measured and predicted collision kernel for non-settling particles are additionally shown as open symbols and broad dotted lines, respectively. The colour scheme in panel (b) follows that of panel (a). (c) The ratio of the actual collision kernel $\Gamma^{sim}$ to the model prediction at $St_{b}=2.8$ and (d) $St_{b}=6.3$ for settling particles. The colour scheme in panel (d) follows that of panel (c).
@@ 5.1 Model validity for the cases with lower bubble Stokes number
where $\rho_{w}(\tau)=\langle w_{b}(t)w_{b}(t+\tau)\rangle_{t}/\langle w_{b}^{2}(t)\rangle_{t}$ is the autocorrelation function of the bubble slip velocity, $\langle\cdots\rangle_{t}$ denotes averaging over time, and $\tau_{\infty}$ is defined as the value of $\tau$ where $\rho_{w}(\tau)$ first decays to 0.01. $T_{L}^{w}/\tau_{f}\gtrsim 1$ would suggest that the frozen turbulence assumption is conceptually valid and the model can be applied. The measured values of $T_{L}^{w}/\tau_{f}$ is shown in figure 7. $T_{L}/\tau_{f}$ generally increases with $1/Fr$ and is far less sensitive to $St_{b}$, such that for all $St_{b}$ considered here the frozen turbulence assumption can be applied when $1/Fr\geq 4$. For $St_{b}=0.5$ and 1, we expect that modelling the slip velocity p.d.f. $f(w_{b})$ by (14) and (26) might slightly underpredict the collision kernel, since the mean value of the slip velocity magnitude is underpredicted as shown in figures 4(ab). This discrepancy in $f(w_{b})$ for $1/Fr\geq 4$ is not due to using the mean vertical rise velocity instead of the magnitude of the bubble rise velocity in the bubble drag correction factor $f_{b}$ in (26), because the bubble velocity at high $1/Fr$ is dominated by its vertical component. As evidence, we plot $f(w_{b})$ with $\sigma_{i}^{(BC)}$ determined using the average $f_{b}$ obtained from the simulations as dashed green lines in figure 4. Only at small $1/Fr$ does using the actual value of $f_{b}$ noticeably improve the quality of the predicted slip velocity p.d.f.. If desired for these cases, a more accurate estimate of $\sigma_{i}^{(BC)}$ can be obtained a priori by solving (26) iteratively using $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ as an initial guess. We additionally note that particles with $St_{p}\sim 1$ preferentially sample flow regions in turbulence which can reduce $\Gamma$ by up to 20% (Jiang and Krug, 2025). Similarly preferential sampling occurs for bubbles with $St_{b}\sim 1$ but they cluster at different regions, which means that bubbles and particles segregate spatially when $(St_{b},St_{p})\sim(1,1)$ unless $1/Fr\gtrsim 10$ where segregation becomes negligible under strong gravity (Chan et al., 2025). This segregation effect is not included in our model and reduces the collision kernel, hence for $(St_{b},St_{p})\sim(1,1)$ at intermediate values of $1/Fr$ the predicted value of the collision kernel may be closer to the actual value than anticipated.
@@ 5.2.1 Relevant parameter space
[…]
@@
5.2.2 Frozen turbulence approach in practice
We now reflect on the applicability of the frozen turbulence approach at practically relevant parameters, which has a higher $Re_{\lambda}$ than the simulated cases. As discussed in §5.1, the frozen turbulence approximation is conceptually valid when $T_{L}^{w}/\tau_{f}\gtrsim 1$. However, estimating the value of $T_{L}^{w}/\tau_{f}$ a priori at a different value of $Re_{\lambda}$ is challenging. When $Re_{\lambda}$ increases, $\langle w_{b}\rangle$ decreases following Eq. (14) so $\tau_{f}$ increases according to Eq. (35). Simultaneously, $\tau_{\eta}$ decreases such that $St_{b}$ increases, which according to our simulated cases increases $T_{L}^{w}$. The net effect of changing $Re_{\lambda}$ on $T_{L}^{w}/\tau_{f}$ is therefore uncertain.
As a crude estimate, we refer to our simulations and note that $T_{L}^{w}/\tau_{f}\gtrsim 1$ holds for $0.5\leq St_{b}\leq 6.3$ when $1/Fr\geq 4$. We furthermore note that the estimated value of $Re_{\lambda}\sim 100$ is not orders of magnitude larger than the simulated value of $Re_{\lambda}=64$. Therefore, we conjecture that for practically relevant conditions, the frozen turbulence approximation holds for $1/Fr\gtrsim O(1)$, which corresponds to $\varepsilon\gtrsim 0.2\,$\mathrm{W}\mathrm{/}\mathrm{k}\mathrm{g}$\$. We emphasise that the above is a very rough estimate and $T_{L}^{w}/\tau_{f}\gtrsim 1$ is a conceptual limit. In practice, the deviation of the model prediction from reality may remain minor even when the conceptual limit is violated. This is exemplified by the $(St_{b},1/Fr)=(2.8,2)$ cases in figures 5(c) and 6(c), where the deviation of $\Gamma^{model}$ from $\Gamma$ is generally less than 30% even though $T_{L}^{w}/\tau_{f}<1$. In view of the uncertainty in the value $\varepsilon$ where the frozen turbulence approximation becomes invalid, we present the model predictions throughout the entire range of $\varepsilon\in[0.1,100]\,$\mathrm{W}\mathrm{/}\mathrm{k}\mathrm{g}$\$.